OpenStudy (anonymous):

could i get some help on how to do this? thanks What is the slope of the line that passes through (-2, 2) and (-4, 5)?

OpenStudy (anonymous):

\[m=(y _{2}-y _{1})/(x _{2}-x _{1})\] Where m = gradient r u familiar with this formulae?

OpenStudy (skullpatrol):

Making a sketch may help.

OpenStudy (anonymous):

well learn it :)

OpenStudy (anonymous):

ill just show u have to use it and ull understand

OpenStudy (anonymous):

m = (5 - 2)/(-4 + 2) m = -3/2 and btw the slope of the graph is the gradient

OpenStudy (anonymous):

(y2 , x2) (y1 , x1) just note that y2 and x2 pair up in the same coordinate

OpenStudy (skullpatrol):